OpenStudy (anonymous):

Triangle ABC has side lengths 7, 24, and 25. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (cwrw238):

if 25^2 = 7^2 + 24^2 then they do 25^2 = 625 7^2 = 49 24^2 = 576 so what do u think?

OpenStudy (cwrw238):

so the answer is yes .